OpenStudy (anonymous):

Factor completely: 18x2 − 21x − 15 3(2x + 1)(3x − 5) 3(2x − 5)(3x + 1) 3(2x − 1)(3x + 5) 3(6x + 1)(x − 5)

OpenStudy (anonymous):

18x^2-21x-15=3(6x^2-7x-5)=3{6x^2-10x+3x-5)

OpenStudy (anonymous):

now can u?

OpenStudy (anonymous):

well it can be written this way as 3(6x^2-7x-5) as w are taking out the common next break the middle term can you do it now?

OpenStudy (anonymous):

WAIT WHAT

OpenStudy (anonymous):

see after taking out the common of the equation you will get 6x^2-7x-5 then you are going to break the middle term ie 7x in such a way that when you multiply yo get 30 and if you add or sub you get 7x so we will get (6x^2-10x+3x-5) now you got it?

OpenStudy (anonymous):

OK YES

OpenStudy (anonymous):

so is it D?